OpenStudy (anonymous):

Find the slope and the y-intercept. 8- ¼ y=2x. The slope is 32. The y intercept is (0, -8), correct?

OpenStudy (anonymous):

1/4 y = -2x + 8 y = -8x + 32 slope is -8 y - intercept is (0, -8)

OpenStudy (anonymous):

oops..atleast i got the y-intercept correct..thank you moneybird..

OpenStudy (anonymous):

sorry y-intercept is (0, 32)

OpenStudy (anonymous):

oh, i got it backwards?

OpenStudy (anonymous):

f(0)=-8*0+32=32

OpenStudy (anonymous):

yeah somehow :D

OpenStudy (anonymous):

ugh, that happens to me all the time..thanks again